# EgoGaussian: Dynamic Scene Understanding from Egocentric Video with 3D Gaussian Splatting

### v1

@@ 3.4 Dynamic object modeling
Figure 3: Object pose estimation pipeline. We use a sequential pipeline with regularization from previous frames, allowing us to simultaneously estimate the object poses and reconstruct the object for each transformation we model, while preserving the correct geometry of the object.
@@
Combining the static scene and dynamic object. As a final step, we combine the object model estimated in Section3.4 with the full background model in Section3.3. In practice, we note that at this stage, there are often floaters belonging to the background that obscure parts of the object. To eliminate these, we perform a final fine-tuning stage using all training frames and Gaussians. As we focus here on optimizing how the background and dynamic object interact and fit with each other, we again freeze the per-frame object pose. This produces then the full scene reconstruction, including per-frame data of the object pose.
@@ 4 Experiments
We compare our method with existing baselines for the dynamic scene reconstruction task. The goal of this task is to reconstruct both static 3D scenes and dynamic objects from RGB egocentric videos.
@@ 4.1 Novel View Synthesis
Datasets. We evaluate our method for dynamic novel view synthesis on in-the-wild videos extracted from the two commonly used egocentric video datasets.
@@
HOI4D [35] is a large-scale egocentric video dataset of human-object interactions, where each video is $20$ seconds long. We randomly select $4$ videos involving active objects undergoing rigid transformations from this dataset. Among these, $2$ videos contain mostly translations, while the other $2$ include both translations and rotations. Compared to the original dataset, we downsample the image resolution to one-quarter of its original size, resulting in a resolution of $480$ x $270$ pixels.
@@
EPIC-KITCHENS [11] is a large-scale dataset featuring in-the-wild egocentric videos of human-object interactions in native kitchen environments. We select $4$ video clips involving rigid transformations from the EPIC-KITCHENS dataset. Of these clips, 2 contain mostly translations, while the other 2 include both translations and rotations. Similar to the HOI4D dataset, we also downsample the image resolution to $455$ x $256$. The average length of these clips is $10.43$ seconds with $60$ FPS.
@@
Evaluation protocol. For each video, we train on every second frame, and evaluate on the rest. Although we are able to correctly track the trajectory of the motion with fewer training frames, the fast motion that comes with egocentric videos leads to a significant loss of information when doing so. As such, we use a much denser concentration of training frames.
@@
Metrics. To assess the performance of our model, we use the peak signal-to-noise ratio (PSNR), the structural similarity index (SSIM)[60], and the VGG-based perceptual similarity metric LPIPS [75]. As we aim to reconstruct the background and object without the actor, we mask out the arm and body of any actors within the scene when computing these metrics, and only evaluate the quality of the object and background reconstruction.
Baselines. We compare our model’s results with the current state-of-the-art (SOTA) methods, Deformable 3DGS [68] and 4DGS [63]. All the compared methods have the publicly available codebase. We can run the code as it is and report the results. All the numbers reported in the tables are benchmarked on a single NVIDIA RTX 6000 or TITAN Xp GPU.
@@
Results. Table 1 and Figure 5 compares our method with existing dynamic Gaussian Splatting methods. We observe that EgoGaussian significantly outperforms existing methods on all evaluation metrics over two datasets. The two SOTA methods have very close performance on both datasets and all methods have better results on static evaluation frames than on dynamic ones.
Table 1: Comparison with SOTA dynamic Gaussian Splatting methods. we evaluate our method and two other SOTA baselines on the HOI4D and EK datasets. The best and second best results are bolded and italicized respectively.
@@
4DGS [63] 0.89 25.45 0.13 0.89 25.25 0.13 0.91 33.25 0.12 0.79 22.77 0.24
Def-3DGS [68] 0.90 25.96 0.11 0.89 25.39 0.12 0.93 33.73 0.11 0.81 23.38 0.22
Ours 0.96 31.52 0.07 0.95 30.29 0.09 0.94 34.22 0.10 0.88 28.30 0.17
@@ 4.2 Dynamic modeling
Figure 4: More qualitative results of the reconstructed dynamic scenes. The center figure shows the motion path, while the four F panels display individual trajectories of the reconstructed frames compared to the corresponding ground truth input frames.
We show in Figure 4 how we can reconstruct the scene including the object trajectory both using the original camera motion, or from a fixed point of view.
4.3 Ablation study
Without full scene fine-tuning. We show the necessity of fine tuning the static background and dynamic object as described in Section 3.4 jointly by comparing how our method performs when the background and object are only trained in isolation without fine-tuning on all frames or on the combined scene. As can be seen in Table 2 without full scene fine tuning, quality drops significantly.
Table 2: Ablation study of full scene fine tuning.
@@ 4.3 Ablation study
Estimate poses with larger time gap. We show that our method is also able to estimate poses with larger time gaps, by training on every 6 frames instead of every 2 frames. As seen in Table 3, although PSNR drops, we are still nonetheless able to produce an accurate reconstruction. Note that the metrics were computed only on the dynamic object itself.
HOI4D Epic-Kitchen
Method Static Dynamic Static Dynamic
5 Conclusion
In this paper, we introduce EgoGaussian, the first method capable of simultaneously reconstructing 3D scenes and dynamically tracking 3D object motion from RGB egocentric input alone. We present a novel Gaussian Splatting framework that leverages the dynamic nature of human activities and distinguishes between dynamic interactions and static captures. This approach allows us to reliably reconstruct the static background and track object motion during the dynamic phase while progressively refining the object shape. Our method significantly outperforms previous SOTA baselines evaluated on two in-the-wild egocentric video datasets. Qualitative results demonstrate the high quality of the reconstructed dynamic scenes.
Limitations and discussion. While our method successfully reconstructs dynamic scenes from egocentric videos, several challenges remain. First, as we constrain objects to be rigid, we cannot model elastic or stretchable objects. Furthermore, although we substantially improve on previous methods, notable image artifacts and floaters still remain. Finally, the iterative process of estimating the object pose significantly increases training time.
Potential negative societal impacts.. EgoGaussian shares most of the societal impacts brought by others 3D modeling methods including potential unsolicited 3D reconstruction of, for example, private properties. With EgoGaussian and its focus on rigid inanimate objects, there is a heightened risk of industrial espionage. However, this risk is mitigated by the requirement that individuals must have access to images of the specific place or product to create a 3D model. This limitation reduces the likelihood of such privacy and security breaches.
GT
@@ 5 Conclusion
4DGS [63]
Def-3DGS [68]
HOI Scene1
HOI Scene2
HOI Scene3
GT
Ours
4DGS [63]
Def-3DGS [68]
HOI Scene4
EK Scene1
EK Scene2
EK Scene3 (Fail)
Figure 5: Qualitative comparison with SOTA. We show reconstructions produced by our method and SOTA baselines (4DGS [63] and Deformable 3DGS [68]) from both HOI4D and EPIC-KITCHENS. Our reconstruction demonstrates more accurate reconstructions while baseline approaches fail to handle dynamic interactions.
Appendix A Appendix / supplemental material
A.1 NeRF-based method experiments
Datasets and evaluation protocol.
In this section we present some supplementary results led with the same dataset as in section 4.1 but with a different division of training and validation frames. For a single given video, We use for training nine out of ten frames from static parts and one out of six frames for dynamic parts.
Metrics and Baseline.
We compare our model to some well known NeRF-based architecture able to model dynamic scene. For it we use Neuraldiff [57] and T-NeRF [47]. We used the official implementation with default parameter for NeuralDiff and the implementation provided by [19] for T-NeRF. For T-NeRF we used the default setting for "nerfies" type dataset and the additional preprocessing was borrowed from [45].
Table 4: Results on dynamic NeRF methods. We evaluate two baselines on the HOI4D and EK datasets.
HOI4D Epic-Kitchen
Method Static Dynamic Static Dynamic
NeuralDiff [57] 0.71 21.15 0.32 0.76 23.98 0.26 0.89 31.06 0.13 0.79 24.85 0.24
T-NeRF [47] 0.84 23.00 0.10 0.78 19.96 0.14 0.80 14.52 0.17 0.75 21.33 0.22
GT
NeuralDiff [57]
T-NeRF [47]
HOI Scene 2
HOI Scene 3
HOI Scene 4
Figure 6: Three different scenes from HOI4D, we show reconstructions produced by NeuralDiff [57] and T-NeRF [19]

### v2

@@ 3.3 Static Reconstruction
Figure 2: Dynamic Object Modeling Pipeline. We use a sequential pipeline with regularization from previous frames, allowing us to estimate object poses and iteratively refine their shapes simultaneously.
Dynamic Object Identification. To identify the set of Gaussians associated with dynamic objects, we extract interaction information from nearby dynamic clips and aim to generate masks for objects that have moved or will move. We automatically generate such masks by selecting a random point from the object mask during interaction within the dynamic clips and using it as a prompt for the Track-Anything model [66]. The initial static reconstruction from the previous step allows us to lift 2D masks to a dense 3D reconstruction of the scene, and conversely, project 3D Gaussian points to 2D. Experimental results show that object masks from just $N$ static frames adjacent to the dynamic clips are sufficient to obtain reliable 3D segmentation of the object. For example, consider a static clip $S$ with $T$ frames immediately preceding a dynamic clip where an object is moved by the camera wearer. We obtain segmentation masks for the object in the last $N$ frames of this static clip, i.e., $\{\mathbf{M}_{\text{obj},T-N},\dots\mathbf{M}_{\text{obj},T}\}$, where a pixel value of $=1$ indicates the targeted object and $=0$ represents the rest of the frame. We set $N=5$ in our experiments.
Similar to Gaussian Grouping [69], though more streamline, an additional trainable parameter of label $l$ is then attached to each Gaussian and initialized to a very small value. This label can then be rendered similar to the RGB value as: $L(\mathbf{x}_{p})=\sum_{i\in\mathcal{N}}l_{i}\alpha_{i}\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right)$. This produces a segmentation $\mathbf{L}_{\text{r}}$, upon which we can apply a binary cross entropy loss using the object masks $\mathbf{M}_{\text{obj}}$:
$\mathcal{L}_{\text{BCE}_{l}}=-\left[\mathbf{M}_{\text{obj}}\cdot\ln\left(\sigma(\mathbf{L}_{\text{r}})\right)+(1-\mathbf{M}_{\text{obj}})\cdot\ln\left(1-\sigma(\mathbf{L}_{\text{r}})\right)\right]$
where $\sigma(X)=\frac{1}{1+e^{-X}}$ is the sigmoid activation function. Training the 3D Gaussians with respect to this loss function while freezing all parameters except for $l$ allows us to separate the Gaussians into object $\mathcal{G}_{\text{obj}}$ and background $\mathcal{G}_{\text{bg}}$ based on thresholding $l$ and such 3D segmentation can be used to render targeted object masks from more viewpoints. More details and graphical illustration are in the Supp.
3.4 Dynamic Reconstruction
Given the reconstructed background $\mathcal{G}_{\text{bg}}$ and the initial object $\mathcal{G}_{\text{obj}}$ using the static clips, we refine the object shapes, track their motion through the dynamic clips and update the background with new information revealed by the movement of the dynamic objects.
Object Pose Estimation. We use the set of object Gaussians $\mathcal{G}_{\text{obj}}$ as an initial estimate of the object appearance and estimate its pose for every $k$-th frame in the dynamic clip. Specifically, we estimate for a targeted frame $f_{t}$ a corresponding relative pose $\mathbf{P}_{t}$ from a previous state at $f_{t-k}$. We further decompose the pose $\mathbf{P}_{t}$ into a 3D translation vector $T_{t}$ and a rotation matrix $\mathbf{R}_{t}$. Unlike previous dynamic 3D-GS methods [63, 36], EgoGaussian applies one set of transformation parameters to the whole collection of object Gaussians $\mathcal{G}_{\text{obj}}$ as a whole, treating it as a single rigid object.
We optimize the rotation $\mathbf{R}_{t}$ using the 6D continuous rotation representation $\tilde{\mathbf{R}}_{t}$ proposed by [77]. To ensure the transformation is rigid, an estimated $3\times 3$ rotation matrix must be $\mathbf{R}_{t}\in SO(3)$. Hence for each target frame $f_{t}$, we apply the estimated translation and rotation parameters to the 3D center of each Gaussian:
$\mathbf{X}_{\mathcal{G}_{\text{obj},t}}=\mathbf{X}_{\mathcal{G}_{\text{obj},t-k}}\cdot g(\tilde{\mathbf{R}}_{t})+\mathbf{t}_{t},$
where $\mathbf{X}_{\mathcal{G}_{\text{obj},t}}$ is the 3D coordinates of object Gaussians $\mathcal{G}_{\text{obj},t}$ at time $t$ and $\mathbf{X}_{\mathcal{G}_{\text{obj},t-k}}$ at time $t-k$. $g(\cdot)$ is a function defined in [77] that transforms the 6D representation of rotation to a standard $3\times 3$ rotation matrix. Such object rotation is also applied to the anisotropic covariance of each 3D Gaussian to regularize its alignment with the object surface:
$\Sigma^{\prime}=\mathbf{R}\Sigma\mathbf{R}^{T}=(\mathbf{R}_{t}\mathcal{R})\mathcal{S}\mathcal{S}^{T}(\mathbf{R}_{t}\mathcal{R})^{T},$
where the covariance matrix $\Sigma$ is decomposed into the scaling vector $\mathcal{S}$ and the Gaussian’s rotation matrix $\mathcal{R}$ constructed from quaternion.
Object Shape Refinement. In order to better reconstruct the shape of the object, we apply a silhouette loss onto the computed alpha value using the following equation: $A(\mathbf{x}_{p})=\sum_{i\in\mathcal{N}}\alpha_{i}\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right)$, which effectively equates the RGB rendering equation without color.
Dynamic Object Reconstruction. Our final dynamic object reconstruction loss is then:
$\mathcal{L}_{\text{obj}}=\mathcal{L}_{1}\left(\mathbf{I}_{\text{obj}},\mathbf{I}_{\text{render},\mathcal{G}_{\text{obj}}}\right)+\lambda\mathcal{L}_{2}\left(\mathbf{M}_{\text{obj}},\mathbf{A}_{\text{render},\mathcal{G}_{\text{obj}}}\right),$
where $\mathbf{I}_{\text{obj}}$ is cropped from $\mathbf{I}_{\text{input}}$ with the object segmentation mask $\mathbf{M}_{\text{obj}}$. $\mathbf{I}_{\text{render},\mathcal{G}_{\text{obj}}}$ is rendered from $\mathcal{G}_{\text{obj}}$ so it contains the object only and black background. $\mathbf{A}_{\text{render},\mathcal{G}_{\text{obj}}}$ is the rendered alpha. We experimentally observe that $0.5$ is a suitable $\lambda$ value. Additionally, we lower the learning rate on the Gaussians parameters, such as position and color, by a factor of 10 to prioritize the learning of object poses.
As we optimize the time-dependent pose parameters $T_{t}$ and $\tilde{\mathbf{R}}_{t}$ one frame at a time, the Gaussians can easily overfit to the current frame. To address this, we train not only on the current frame, instead for every training iteration we train either on the current frame or a random previous frame with a probability of $0.5$.
Static Scene Update. The static scene $\mathcal{G}_{\text{bg}}$ is reconstructed from the static clips, where parts of the background are obscured by objects that interact within the dynamic clips. To utilize the visible information from the dynamic clips, we retrain $\mathcal{G}_{\text{bg}}$ with the dynamic objects masked out.
@@ 3.4 Dynamic Reconstruction
Combination of Static Scene and Dynamic Objects. As a final step, we combine the object model $\mathcal{G}_{\text{obj}}$ with the full background model $\mathcal{G}_{\text{bg}}$. In practice, we note that at this stage, there are often floaters belonging to the background that obscure parts of the object. To eliminate these, we perform a final fine-tuning stage using all training frames and Gaussians. As we focus here on optimizing how the background and dynamic object interact and fit with each other, we again freeze the estimated per-frame object pose. This produces then the full scene reconstruction, including per-frame data of the object pose.
Training details. During the Gaussian splatting optimization process, the opacity is frequently set to zero in order to prune floaters. However, this would produce a very noisy signal for the pose optimization. As such, we instead alternate between optimizing the rigid object pose, and densifying/pruning the Gaussians. We first train for 4k iterations on every dynamic frame, optimizing the object poses without pruning or densification. We then freeze the object poses, and train another 4k iterations to better incorporate visual information with the estimated object poses, and finally train another 4k iterations optimizing the poses without densification or pruning. For all 12k iterations, all Gaussian parameters such as color and position are continuously optimized. After iterating through the whole dynamic clip with $M$ frames, we obtain a coarse object pose for each frame $P=\{T_{t},\mathbf{R}_{t}|t=1,\ldots,M\}$. Finally, we perform one final round of joint training using all frames, with 6k iterations of pose estimation, 6k iterations of pruning/densifications, and finally another 6k iterations of pose estimation. This ensures that our object model is more equally fit onto all frames, rather than focused on the last seen ones.
@@ 4 Experiments
We compare our method with existing baselines for dynamic scene reconstruction where the goal is to reconstruct both static 3D scenes and dynamic objects from RGB egocentric videos. To quantitatively assess the quality of the reconstructed 4D scene, we follow the evaluation protocol of the novel view synthesis task using two different egocentric video datasets. We then present qualitative results of the reconstructed dynamics and conduct ablation studies on two key aspects of the proposed method.
@@ 4.1 Novel View Synthesis
Datasets. We evaluate our method on two commonly used egocentric video datasets HOI4D and EPIC-KITCHENS.
@@
HOI4D [33] is a large-scale egocentric video dataset of human-object interactions consisting of $20$ second-long videos. From this dataset, we randomly select $4$ videos involving rigidly moving objects. Among these, $2$ videos contain mostly translations, while the other $2$ include both translations and rotations. Compared to the original dataset, we downsample the image resolution to one-quarter of its original size, resulting in a resolution of $480$ x $270$ pixels. Each video has a framerate of $15$ FPS.
@@
EPIC-KITCHENS [11] is a large-scale dataset featuring in-the-wild egocentric videos of human-object interactions in native kitchen environments. We again randomly select $4$ video clips involving rigidly moving objects. Of these clips, 2 contain mostly translations, while the other 2 include both translations and rotations. Similar to the HOI4D dataset, we also downsample the image resolution to $455$ x $256$. The average length of these clips is $10.43$ seconds with $60$ FPS.
@@
Evaluation protocol. For each video, we train using every second frame and evaluate on the rest. Although we are still able to correctly track and even reconstruct the moving object with fewer training frames (i.e. frames with larger step size), the rapid motion that comes with egocentric videos means that neither our method nor any baselines are able to properly reconstruct the background. We show examples with different step sizes in the Supp.
@@
Metrics. To assess the performance of our model, we use the peak signal-to-noise ratio (PSNR), the structural similarity index (SSIM)[60], and the VGG-based perceptual similarity metric LPIPS [75]. As we aim to reconstruct the background and object without the actor, we mask out the arm and body of any actors within the scene when computing these metrics and only evaluate the quality of the object and background reconstruction.
Baselines. We compare our method with two state-of-the-art dynamic 3D-GS-based methods, Deformable 3DGS [68] and 4DGS [63], where both methods apply deformation fields to model monocular dynamic scenes. All compared methods have publicly available codebases, and as such we run the code as provided by the authors on the same preprocessed egocentric data used by our method.
However, unlike our method, both of the SOTA approaches are unable to properly utilize masks of image regions which should not be modeled, in this case the hands of the actor. To ensure a fairer comparison, we additionally modify the compared methods to support masking out gradients on the segmented human body, similar to our method.
@@
Results. Table 1 and Figure 3 compares our method with existing dynamic 3D-GS methods and their modified versions. We observe that, though the adaptation of gradient masks increases the performance of existing methods, EgoGaussian still significantly outperforms both methods on all evaluation metrics over the dynamic frames of two datasets, with the two SOTA methods performing similarly on both datasets.
The supplementary video shows the sequences from which the images originate separately. The video results suggest that the deformation-field-based method is not able to track the rigid movement of objects in egocentric video and tends to overfit to the training views.
@@
4DGS [63] 0.88 25.33 0.13 0.89 25.34 0.13 0.84 26.84 0.20 0.79 22.54 0.24
4DGS w/o hands 0.94 28.69 0.08 0.94 27.33 0.10 0.87 28.90 0.16 0.80 23.13 0.23
Def-3DGS [68] 0.90 25.85 0.11 0.90 25.71 0.12 0.86 27.35 0.18 0.81 23.15 0.22
Def-3DGS w/o hands 0.94 28.09 0.08 0.94 26.92 0.10 0.86 27.63 0.17 0.82 23.27 0.21
Ours 0.96 30.99 0.08 0.95 30.33 0.09 0.85 28.33 0.19 0.88 28.34 0.17
Table 1: Comparison with SOTA dynamic Gaussian Splatting methods. We evaluate our method and two other SOTA baselines along with their modified versions with hands excluded from modeling on the HOI4D and EK datasets. The best and second best results are bolded and italicized respectively. We show the evaluation results on static and dynamic frames separately.
GT
Ours
4DGS [63]
4DGS w/o hands
Def-3DGS [68]
Def-3DGS w/o hands
HOI Scene1
HOI Scene2
HOI Scene3
HOI Scene4
EK Scene1
EK Scene2
Figure 3: Qualitative comparison with SOTA. We show reconstructions produced by our method and SOTA baselines (4DGS [63] and Deformable 3DGS [68]) along with their modified versions from both HOI4D and EPIC-KITCHENS. Our reconstruction demonstrates more accurate reconstructions while baseline approaches fail to handle dynamic interactions.
@@ 4.2 Dynamic modeling
Figure 4 shows the estimated object trajectories and novel views rendered from arbitrary viewpoints. Demonstration videos and more visualizations are included in the Supp.
Figure 4: More qualitative results of the reconstructed dynamic scenes. The left figure shows the object trajectory inferred from interpolated object poses. The right figure shows the rendering from arbitrary viewpoints.
HOI4D Epic-Kitchen
Method Static Dynamic Static Dynamic
@@
Table 3: Ablation study of full scene fine tuning.
4.3 Ablation study
Figure 5: Example of black artifacts
@@ 4.3 Ablation study
Estimate poses with larger time gap. We show that our chronological pose estimation schema is also able to model the dynamic object with larger time gaps $k$, by training on every 6 frames instead of every 2 frames. We can then estimate the state of object at each timestamp $t$ through interpolation of transformation matrix constructed from relative pose $\mathbf{P}_{t}$ from $t-k$ to $t$. As seen in Table 2, although PSNR drops, we are still nonetheless able to produce an accurate reconstruction. Note that the metrics are computed only on the dynamic object itself.
Without full scene fine-tuning. We show the necessity of fine-tuning the static background and dynamic object as described in Section 3.4 jointly by comparing how our method performs when the background and object are only trained in isolation without fine-tuning on all frames or on the combined scene. As can be seen in Table 3 without full scene fine-tuning, quality drops significantly. This is partially due to 3D-GS being unable to distinguish between transparency and blackness. By jointly training both object and background, we eliminate this uncertainty and such black artifacts. We show examples of this in Figure 5.
5 Conclusion and Discussion
We introduced EgoGaussian, a novel egocentric reconstruction method that is able to reconstruct rigid objects along with accompanying motion from egocentric data. We show significant improvements in terms of both dynamic object and background reconstruction quality compared to the state-of-the-art.
Although our method is able to well reconstruct rapid, rigid object motion, there are still a number of important limitations. As we require labels for both onset and offset of objects motion, as well as object masks, we require several additional offline data-preprocessing steps. Our method fundamentally relies on multiview-stereo data in order to reconstruct the scene geometry. As such, our method can encounter overfitting if the datasets have limited viewpoint coverage or lack features that can be tracked across time or view (e.g. uniformly textured walls). We utilize pixel-wise gradients to estimate per-frame object motion. As such, if the object goes out of frame, or there is too much motion between two frames, our method can lose track of the object entirely. The frame-by-frame dynamic object modelling requires significantly longer training compared to most previous dynamic 3D-GS-based methods.
Both motion labels and mask segmentations are parallel avenues of ongoing research, and as improvements are made, can easily be integrated into our pipeline. Similarly, as our method fundamentally trains using static or rigid sets of Gaussians, future improvements in sparse 3D Gaussian reconstruction could also be integrated into our pipeline. Although pixel-wise gradients prove to be sufficient in our setting, non-local methods of supervision such as optical flow would be an interesting avenue of research in order to improve robustness for large motions or when objects move out of frame. Such forms of supervision could also be utilized in order to estimate per-frame object pose more rapidly, allowing for reduced training time.
Supplementary Material
6 Static Reconstruction Pipeline
A graphical illustration of our static reconstruction pipeline is provided in Figure 6. For each static clip from the egocentric video, we get two Gaussian sets for static background and dynamic object respectively, where the masks projected from object Gaussians are used to complete the previously occluded parts of background.
7 More Experiment Results
7.1 Reconstruction of Dynamic Object
We also compare our methods with Deformable 3DGS [35] and 4DGS [63] on datasets of frames that exclusively contain the targeted object to exclude the effects from both static background and human body. The datasets are preprocessed to crop out everything except the dynamic object; gradient updates are also disabled on hands that may obscure the object during interaction.
As illustrated in the supplementary video, the deformation-field-based methods encounter difficulties tracking and modeling the dynamic object–Def-3DGS [68] can only model the very initial movement while 4DGS [63] does not yield any meaningful rendering. Our method is able to accurately track and therefore reconstruct the object with the sequential pose estimation.
7.2 Pose Estimation with Larger Time Gap
In Table 2, we show that we are still able to model the dynamic objects in the scene accurately with a larger time gap in pose estimation. This exchanges a slight performance cost for a significant reduction in training time. We further demonstrate this qualitatively in video results and visualized object trajectories. For simple movements where translation is dominant, the object reconstruction quality and estimated trajectory after pose interpolation remain similar; while for more complicated movements involving rotation, we observe degradation in reconstruction quality
Although our method can track the dynamic object even across larger time gaps, we cannot properly reconstruct the background, likely due to much sparser information in the presence of strong camera motion. This applies not only to our method but also to both 4DGS and Deformable 3DGS. See examples in Figure 7.
7.3 Failure Cases
We mainly observe two types of reconstruction failure. The first type is when the interacted object moves out of sight, which sometimes happens in egocentric video for cameras with a wider field of view, the pose estimation can get stuck in a local minima and lose track of the object in the subsequent frames. The second type arises when the camera registration quality is poor, which is often due to featureless surfaces and insufficient overlap between video frames. The inaccurate camera poses result in a static reconstruction that lacks any geometric understanding of the scene, thereby causing the dynamic reconstruction to fail.
Figure 6: Static Reconstruction Pipeline. We use this pipeline to jointly reconstruct and segment dynamic object from static background from a static egocentric video clip.
@@ 7.3 Failure Cases
4DGS
Def-3DGS
Figure 7: Comparison with SOTA methods on background reconstruction with larger time gap All methods suffer from floaters due to sparse training viewpoints